### v1

# Learning Quadruped Locomotion using Differentiable Simulation

Abstract
While most recent advancements in legged robot control have been driven by model-free reinforcement learning, we explore the potential of differentiable simulation. Differentiable simulation promises faster convergence and more stable training by computing low-variant first-order gradients using the robot model, but so far, its use for legged robot control has remained limited to simulation. The main challenge with differentiable simulation lies in the complex optimization landscape of robotic tasks due to discontinuities in contact-rich environments, e.g., quadruped locomotion. This work proposes a new, differentiable simulation framework to overcome these challenges. The key idea involves decoupling the complex whole-body simulation, which may exhibit discontinuities due to contact, into two separate continuous domains. Subsequently, we align the robot state resulting from the simplified model with a more precise, non-differentiable simulator to maintain sufficient simulation accuracy. Our framework enables learning quadruped walking in minutes using a single simulated robot without any parallelization. When augmented with GPU parallelization, our approach allows the quadruped robot to master diverse locomotion skills, including trot, pace, bound, and gallop, on challenging terrains in minutes. Additionally, our policy achieves robust locomotion performance in the real world zero-shot. To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot. This work provides several important insights into using differentiable simulations for legged locomotion in the real world.
I Introduction
Many grazing animals, such as giraffes, lambs, and calves, can learn to stand and walk within minutes to hours after birth. On the other hand, most machine learning algorithms, particularly traditional model-free reinforcement learning (RL), need massive parallelization to achieve a stable walking policy. How can legged robots master walking in minutes with a few trials and errors?
Recent progress in legged robot control has largely been driven by GPU-accelerated massive simulation and reinforcement learning [15, 20, 24, 30, 22, 9, 39, 19]. Central to these advancements is the policy gradient algorithm [35, 34, 6]. The policy gradient can generally be expressed as the following
$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R(\bm{\tau})],$
where it stands as a zero-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$. Despite impressive real-world performance, this approach to gradient estimation is known for its high variance. Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [31], increasing the number of samples, designing task-specific learning curriculums, employing distributed initialization techniques, and engineering suitable reward functions.
As a result, RL struggles with complex tasks where data generalization is slow, like vision-based navigation. Despite the need for more effective learning algorithms, progress has been limited. Instead, research leans towards imitation learning, utilizing existing datasets or expert demonstrations for more consistent learning outcomes. Notably, the learning-by-cheating approach [4] is a popular imitation learning framework: first, you train an RL teacher with privileged information and then, you derive a student policy based on visual inputs. However, the success of this approach heavily depends on the quality of expert demos and lacks generalizability and adaptability to new tasks.
In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zero-order counterparts and hold great potential for more stable training and faster convergence. Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zero-order methods.
Figure 1: Graphical model for policy learning using differentiable simulation, in which both the loss function and the dynamics are differentiable.
However, the optimization landscape in robotic tasks is inherently complex, characterized by nonlinearities, non-smooth and discontinuous dynamics, non-convexities, and long temporal horizons. Additionally, deploying real-world robots introduces additional challenges, such as the sim-to-real gap, noisy state estimation, and system delay. Consequently, policy training through differentiable simulation can suffer from issues like local minimum and gradient vanishing/explosion. While differentiable simulation offers a promising avenue, its effectiveness in significantly accelerating policy learning in contact-rich environments and its applicability to real-world scenarios like dynamic quadruped control remains an open challenge in robotics.
This work set out to explore the potential of differentiable simulation in accelerating policy training within contact-rich scenarios like dynamic legged locomotion. Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes. Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zero-order methods, particularly model-free policy gradients. We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
Contribution: We show that differentiable simulation offers considerable advantages over model-free reinforcement learning for policy training in legged locomotion. Notably, we demonstrate that a single robot, without the need for parallelization, can quickly learn to walk within minutes using our approach. Leveraging the advantage of GPUs for parallelized simulation, our robot learns diverse walking skills over challenging terrain in minutes. Specifically, we train a quadruped robot to walk with different gait patterns, including trot, pace, bound, and gallop, and with varying gait frequencies. While model-free RL can also achieve similar performance given sufficient parallelization, our approach requires much less data and yields better performance.
More importantly, we show that the policy trained via differentiable simulation can be transferred to the real world directly without fine-tuning. To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot. This highlights the effectiveness of differentiable simulation for real-world applications in dynamic legged locomotion scenarios.
The key to our approach is a novel policy training framework that combines the smooth gradients obtained from a simplified dynamics model for efficient backpropagation with the high fidelity of a more complex, non-differentiable simulator for accurate forward simulation. 1) Decoupling Simulation Spaces: Rather than handling the whole body dynamics as a unified differentiable simulator, which may encounter non-smoothness due to contact interactions. We proposed to separate the simulation into two distinct spaces: the floating base space and the joint space. For the robot body, we employ an approximation using single rigid-body dynamics, which offers a continuous and effective representation of the floating base (the robot body). 2) PD Control as A Differentiable Layer: The simulation within the joint space benefits from the inherent differentiability and smoothness of a Proportional-Derivative (PD) controller. This allows us to treat the PD control as an explicit, differentiable layer within our simulation framework. 3) Alignment with Non-Differentiable Simulators: To address potential inaccuracies arising from our simplified rigid-body dynamics model, we incorporate a more precise, non-differentiable simulator. This non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our training pipeline remains grounded in realistic dynamics.
II Related Work
II-A Model Predictive Control
Model predictive control (MPC) has a long history in legged locomotion [18, 16]. MPC relies on fast online optimization and an accurate dynamics model to control the robot. To trade-off between model accuracy and computation speed, many methods use a reduced optimization model, such as single rigid-body dynamics [7, 11, 2]. Research has also studied control using more complex formulations, including centroidal [10] and whole-body dynamics [26]. Despite their success, MPC lacks the flexibility in its optimization objective design and robustness when facing disturbances or unexpected scenarios. Empirical evidence suggests that MPC tends to underperform when compared to data-driven approaches such as deep reinforcement learning.
II-B Reinforcement Learning
Deep reinforcement learning has emerged as a dominant approach for developing control policies in legged locomotion [15, 20, 24, 19]. Recently, an important advancement in this area has been the introduction of IsaacGym [21], which is a GPU-accelerated simulator. IsaacGym can dramatically reduce the time required for dynamics simulation and policy training, enabling robots to learn to walk on flat ground in minutes [30]. This advancement has significantly accelerated the pace of research in legged locomotion, opening new avenues to address increasingly complex challenges [38, 1, 9, 22, 39].
Nevertheless, it is important to highlight that the progress in applying reinforcement learning to legged locomotion is primarily driven by the enhanced computational capabilities provided by modern GPUs rather than substantial breakthroughs in the underlying algorithms. Consequently, in scenarios where data collection cannot be accelerated through computational means, researchers may resort to alternative strategies such as “learning-by-cheating” to circumvent these limitations. This highlights a gap in the current approach to deep RL in legged locomotion, suggesting a need for innovative algorithmic developments that do not solely rely on computational advancements.
II-C Differentiable Simulation
Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27]. In principle, policy training through differentiable simulation allows better convergence by replacing the zero-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33]. However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
Several approaches have been developed to tackle these challenges. For instance, SHAC [36] addresses the gradient issue by truncating trajectories into several smaller segments, which helps manage exploding or vanishing gradients. On the other hand, PODS [25] takes advantage of second-order gradients. This approach facilitates monotonic policy improvement and promises faster convergence than first-order methods. However, PODS’ effectiveness relies on very accurate computation of second-order derivatives from the differentiable simulator, making it sensitive to the precision of these derivatives. Another notable example is DiffMimic [29], which focuses on mimicking motion trajectories for physically simulated characters.
Research in differentiable simulation has mainly been focused on speeding up policy optimization using first-order gradients, alleviating local minima problems, addressing gradient vanishing/explosion issues, or developing control policies for simulated characters. They show that differentiable simulation is a promising direction for improving sample efficiency. However, the capability of differentiable simulation to substantially speed up policy learning in contact-rich scenarios with physical interactions, as well as their practical applicability to real-world situations, such as quadruped locomotion, continues to be a significant challenge in robotics.
Our research distinguishes itself from earlier studies by presenting a novel perspective that leads to an effective solution tailored for real-world robotic applications.
III Methodology
Figure 2: System overview of legged robot control using differentiable simulation. Our approach leverages the differentiability and smoothness of a single rigid-body dynamics for dynamics simulation. Additionally, we treat PD control as a differential layer in our computation graph. Finally, we use the state from a more accurate, non-differentiable simulator (IsaacGym) to align the simulated state in the simple rigid-body simulation.
III-A Overview
Fig. 2 shows our training setup that employs differentiable simulation. The fundamental idea is that the main body of a legged robot can be modeled using differentiable and smooth single rigid-body dynamics, while the simulation of the robot’s four legs can be approached separately from its main body. We use a Proportional-Derivative (PD) controller to connect the legs dynamics with the main body. In the forward pass, this PD controller converts joint angles into torques, which then produce the contact forces needed for the robot’s rigid-body simulation. During backpropagation, the PD controller functions as a differentiable component, enabling gradients to be propagated back through the system. Finally, we use non-differentiable whole-body dynamics simulation to align the state produced by our simplified rigid-body dynamics.
III-B Problem Formulation
We formulate legged robot control as an optimization problem. The robot is modeled as a discrete-time dynamical system, characterized by continuous state and control input spaces, denoted as $\mathcal{X}$ and $\mathcal{U}$, respectively. At each time step $k$, the system state is $x_{k}\in\mathcal{X}$, and the corresponding control input is $u_{k}\in\mathcal{U}$. An observation $o_{k}\in\mathcal{O}$ is generated at each time step based on the current state $x_{k}$ through a sensor model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{k}=h(x_{k})$. The system’s dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system as $x_{k+1}=x_{k}+\Delta t\cdot f(x_{k},u_{k})$. The discrete-time instants are $t_{k}=\Delta t\cdot k$, where $k$ ranges from 0 to $N$, thereby establishing a finite time horizon for the control problem. At each time step $k$, the robot receives a cost signal $l_{k}=l(x_{k},u_{k})$, which is a function of the current state $x_{k}$ and the control input $u_{k}$.
The control policy is represented as a deterministic, differentiable function such as a neural network $u_{k}=\pi_{\theta}(o_{k})$. The neural network takes the observation $o_{k}$ as input and outputs the control input $u_{k}$. The optimization objective is to find the optimal policy parameters $\theta^{\ast}$ by minimizing the total loss via gradient descent:
$\min_{\theta}\mathcal{L}_{\theta}$ $=\sum_{k=0}^{N-1}l(x_{k},u_{k})=\sum_{k=0}^{N-1}l(x_{k},\pi_{\theta}(o_{k}))$ (1)
$\theta$ $\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta}$ (2)
where $\alpha$ is the learning rate and $l(x_{k},u_{k})$ is the differentiable loss at simulation time step $k$.
III-C Forward Simulation
Legged robots are characterized by complex dynamics resulting from their multi-body configuration and the contact between their body and the environment. The contacts introduce discontinuities in the dynamics, making first-order gradient-based optimization algorithms difficult. An important aspect of our approach involves separating the simulation of the robot’s body from that of its joints. This design choice allows a simple implementation of the differentiable simulator.
First, we represent the robot’s main body using single rigid-body dynamics. Single rigid-body dynamics have been shown to be useful for dynamic locomotion using online gradient-based optimization [7, 11]. The single rigid-body model offers a continuous representation of the robot base dynamics, avoiding the complex optimization landscape introduced by contacts. We develop our differentiable simulation using the single rigid-body dynamics, which is expressed as follows,
$\dot{\mathbf{p}}_{WB}$ $=\mathbf{v}_{WB}$ $\dot{\mathbf{v}}_{WB}$ $=\frac{1}{m}\sum_{i}\mathbf{f}_{i}+\mathbf{g}$
$\dot{\mathbf{q}}_{WB}$ $=\frac{1}{2}\Lambda(\bm{\omega}_{B})\cdot\mathbf{q}_{WB}$ $\dot{\bm{\omega}}_{B}$ $=\mathbf{I}^{-1}\left(\bm{\eta}-\bm{\omega}_{B}\times(\mathbf{I}\bm{\omega}_{B})\right)$
In this approximation, the state of our system is
$\mathbf{x}=[\mathbf{p},\mathbf{q},\mathbf{v},\bm{\omega}],$
where $\mathbf{p}_{\text{WB}}\in\mathbb{R}^{3}$ is the position and $\mathbf{v}_{\text{WB}}\in\mathbb{R}^{3}$ is the linear velocity of the center of mass in the world frame ${W}$, We use a unit quaternion $\mathbf{q}_{\text{WB}}$ to represent the orientation of the body in the world frame and use $\bm{\omega}_{\text{B}}$ to denote the body rates in the body frame ${B}$. Here, $\mathbf{\Lambda}(\bm{\omega}_{B})$ is a skew-symmetric matrix. The control inputs are the ground reaction force $\mathbf{f}_{i}$ from the legs.
To simulate the rigid-body dynamics, it is important to know the ground reaction force. One option for the control policy design is to output the ground reaction force directly, similar to the MPC design [7, 11]. Another option is to control the robot in the joint space, e.g., output desired joint position, which allows more control authority for the policy and adaptive behavior. In this case, it is required to convert the neural network output (joint position) into the control input of the single rigid-body model, which is the ground reaction force. This conversion is achieved using a PD controller for forward propagation,
$\bm{\tau}=\mathbf{k_{p}(q^{\text{ref}}-q)}+\mathbf{k_{d}(\dot{q}^{\text{ref}}-\dot{q})}$ (3)
which calculates the required motor torque $\bm{\tau}$. Here, $\mathbf{k_{p}}$ and $\mathbf{k_{d}}$ are fixed gains, $\mathbf{q^{\text{ref}}}$ and $\mathbf{q}$ are the reference joint position and the current joint position separately, and $\mathbf{\dot{q}^{\text{ref}}}$ and $\mathbf{\dot{q}}$ are the reference joint velocity and the current joint velocity respectively.
Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$:
$\mathbf{f}=\mathbf{J}\bm{\tau}.$
The continuous nature of PD control enables the backpropagation of policy gradients. Consequently, our simulation framework treats the PD controller as an explicit, differentiable layer.
III-D Backpropagation Through Time
In differentiable simulation for policy learning, the backward pass is crucial for computing the analytic gradient of the objective function with respect to the policy parameters. Following [23], the policy gradient can be expressed as follows
$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{i}}{\partial x_{i}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
where the matrix of partial derivatives $\partial x_{j}/\partial x_{j-1}$ is the Jacobian of the dynamical system $f$. Therefore, we can compute the policy gradient directly by backpropagating through the differentiable physics model and a loss function $l_{k}$ that is differentiable with respect to the system state and control inputs. A graphical model for gradient backpropagation in policy learning using differentiable simulation is given in Fig. 1. Due to the usage of multiplication $\prod$, there are two potential issues in using Eq. (4) for policy gradient: 1) gradient vanishing or exploding, 2) long computation time. We tackle these two problems via short-horizon policy training.
III-E Short-Horizon Policy Training
Although smoothed physical models (e.g., single rigid-body dynamics) improve the local optimization landscape, the complexity of the optimization problem escalates significantly in long-horizon problems involving extensive concatenation of simulation steps. The situation further deteriorates when the actions within each step are interconnected through a nonlinear and nonconvex neural network control policy. The complexity of the resulting reward landscape leads simple gradient-based methods to become trapped in local optima quickly. Instead of directly solving a long-horizon policy training problem, following [36, 30], where long-horizon simulation tasks are truncated into short-horizon simulation, we utilize short-horizon policy learning to achieve stable gradient backpropagation over a short horizon.
III-F Alignment with A Non-Differentiable Simulator
Due to the simplification of our single rigid-body dynamics, the robot state can diverge from its actual one, eventually leading to unrealistic states due to compounding errors. We proposed to align the body state in our differentiable simulation using information from other simulators that use accurate whole-body dynamics. Specifically, we align the robot state in our differentiable simulation with the state information from IsaacGym [21].
We use the following equation to align the robot state.
$\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*(\mathbf{x}^{\text{diff}}_{t+1}-\mathbf{x}^{\text{diff, detach}}_{t+1})$ (5)
Here, $\mathbf{x}^{\text{diff, detach}}_{t+1}$ and $\mathbf{x}^{\text{diff}}_{t+1}$ represent the same robot state of our differentiable simulator at time step $t+1$; hence, they share the same value. The word detach indicates that $\mathbf{x}^{\text{diff, detach}}_{t+1}$ is detached from the computational graph for automatic differentiation. Therefore, we can reset the robot state using the value from the non-differentiable simulation during forward simulation: $\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*\bm{0}$. During backpropagation, the gradient of the state at a given time $t$ is computed as follows
$\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}=\bm{0}+\alpha*\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}.$ (6)
Here, $0<\alpha\leq 1$ is used to decay the gradient.
The non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our differentiable training pipeline remains grounded in realistic dynamics. At the same time, our approach benefits from the simplified differentiable simulator, which offers smooth gradients for backpropagation. Fig. 2 shows the computational graph of the forward propagation and backpropagation using alignment.
III-G Differentiable Loss Function
A fundamental difference between policy training using differentiable simulation and reinforcement learning is that the loss function has to be differentiable when using differentiable simulation. On the one hand, RL allows the direct optimization of non-differentiable rewards, such as binary rewards $0/1$. On the other hand, differentiable simulation requires a smooth differentiable cost function to provide learning signals for the desired control inputs.
We formulate a loss function tailored for velocity tracking, where the main objective is to follow a specified velocity command, denoted as $\mathbf{v}^{\text{ref}}$. Additionally, we maintain the robot’s body height, represented by $p_{z}$. To enhance the robustness of this system, we incorporate several regularization terms: one to mitigate large angular velocity, thus controlling the body rate; another to limit the output action, preventing large control actions; and a term aimed at stabilizing the robot’s orientation using projected gravity vector.
$l(\mathbf{x}_{t},\mathbf{u}_{t})$ $=$ $b\|\mathbf{v}-\mathbf{v}^{\text{ref}}\|^{2}+a\|p_{z}-p_{z}^{\text{ref}}\|+c\|\bm{\omega}\|^{2}$ (7)
$+d\|\mathbf{u}\|^{2}+f\|\mathbf{g}_{\text{proj}}\|+g\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2}$
Finally, we use a foot position loss function to provide a learning signal for the swing legs. This loss is critical for the swing leg since it contains information about the motor position. The swing trajectory is computed by fitting a quadratic polynomial over the lift-off, mid-air, and landing position of each foot, where the lift-off position is the foot location at the beginning of the swing phase, the landing position is calculated using the Raibert Heuristics [28].
IV Experimental Setup
[…]
Observation and Action: The policy observation includes random commands ($\textbf{cmd}_{\text{rand}}$) for the reference velocity, sinusoidal and cosinusoidal representations of gait phases, the base velocity ($\mathbf{v}_{WB}$), the base orientation ($\mathbf{q}_{WB}$), the angular velocity ($\bm{\omega}_{B}$), motor position deviations from default ($\mathbf{q}-\mathbf{q}_{\text{default}}$), and a projected gravity vector ($\mathbf{g}_{\text{projected}}$). The policy action $\delta\mathbf{q}$ is the desired joint position offset from the default joint position.
Observation Dimension Action Dimension
$\textbf{cmd}_{\text{rand}}$ 3 $\delta\mathbf{q}$ 12
$\bm{\sin}(\text{gait phase})$ 4
$\bm{\cos}(\text{gait phase})$ 4
$\mathbf{v}_{WB}$ 3
$\mathbf{q}_{WB}$ 4
$\bm{\omega}_{B}$ 3
$\mathbf{q}-\mathbf{q}_{\text{default}}$ 12
$\mathbf{g}_{\text{projected}}$ 3
TABLE I: Policy observation and action.
[…]
V Experimental Results
V-A Learning to Walk with One Robot
This section explores whether a robot can learn to walk in a single simulation environment without parallelization. We design a simple velocity tracking task where the robot is required to follow a constant velocity in the $x$-axis, e.g., $v_{x}=0.2~{}m/s$. At each training iteration, we simulate 24 times steps with one single robot. Hence, each training iteration contains only 24 data points. We trained the policy for a total 1000 iterations, which corresponded to 24,000 data points and 4 minutes of real-time experiences in total. The result is given by the learning curve in Fig. 3. Surprisingly, despite extremely limited data, our policy successfully learns to walk after minutes of training. As a comparison, model-free reinforcement learning failed to yield positive outcomes under the same condition.
Figure 3: Learning to walk with one simulated robot. We run 10 experiments with different random seeds. The plot is smoothed using a moving average.
V-B Learning Diverse Walking Skills on Challenging Terrains
[…]
Fig. 5 shows the learning curves using different numbers of robots for policy training. We compare the learning performance of our method with a model-free RL algorithm (PPO) [31]. Given minimal samples, e.g., only four robots, RL has a plodding convergence speed and does not learn meaningful walking skills, e.g., the policy constantly falls after a few simulation steps. In contrast, differentiable simulation achieves much higher rewards and can acquire useful walking skills, albeit with relatively low success rates.
As the number of robots increases, the performance for both algorithms improved. Notably, the performance improvement for RL is much more significant than our approach. This indicates the that the zero-order gradient estimates used by reinforcement learning is generally inaccurate and requires much more samples to achieve stable training. On the contrary, the first-order gradient estimates used by our differentiable simulation can have very stable and accurate gradients even given very limited simulation samples. Fig. 5 demonstrates diverse walking skills over challenging terrains using a blind policy trained via differentiable simulation.
Figure 4: Learning to walk on challenging terrains, reinforcement learning versus differentiable simulation. Differentiable simulation achieves faster and more stable convergence, given limited simulation samples.
Additionally, we compare the training wall-clock time and the performance of the resulting policies between the two methodologies. Differentiable simulation often involves the creation of a complex computational graph to facilitate gradient computation, especially for backpropagation through time. As the simulation horizon extends, the length of the computational graph can increase proportionally, leading to a substantial increase in the total training time. This increase can diminish its benefits in situations where gathering a large number of samples is both cheap and fast.
Robots Final Reward Training Time [min]
PPO Ours PPO Ours
4 $-25.82\pm 2.16$ $\mathbf{-3.83\pm 3.01}$ $\mathbf{7.27\pm 0.12}$ $7.28\pm 0.12$
16 $-8.07\pm 1.30$ $\mathbf{-2.03\pm 0.69}$ $7.54\pm 0.06$ $\mathbf{7.50\pm 0.08}$
64 $-6.42\pm 0.29$ $\mathbf{-2.49\pm 0.37}$ $8.46\pm 0.16$ $\mathbf{8.23\pm 0.07}$
1024 $-2.80\pm 0.13$ $\mathbf{-2.69\pm 0.06}$ $\mathbf{11.44\pm 0.13}$ $11.55\pm 0.09$
TABLE II: Learning to walk on challenging terrains, reinforcement learning versus differentiable control. We compare the final reward and the total training walk-clock time.
Figure 5: Diverse locomotion skills over challenging terrains using a blind policy trained via differentiable simulation.
As presented in Table II, a differentiable simulation achieves comparable overall training time. This efficiency is achieved by employing a truncated simulation timestep, which results in a smaller computational graph and enables fast and effective backpropagation. RL requires marginally longer training time due to its mini-batch and multi-epoch training scheme, in contrast, differentiable simulation only uses one gradient update. Differentiable simulation achieved overall higher reward during evaluation.
V-C On the Importance of Non-differentiable Terminal Penalty
This subsection highlights one important benefit of Reinforcement Learning (RL) compared to differentiable simulation: RL can significantly enhance its robustness by directly optimizing through non-differentiable rewards or penalties. Specifically, we use an non-differentiable value $p=200$ to penalize the robot (Eq. 7) when the robot experiences termination during training, e.g., falling on the ground.
$r(\mathbf{x}_{t},\mathbf{u}_{t})=\begin{cases}-l(\mathbf{x}_{t},\mathbf{u}_{t})-p&\text{if termination}\\ -l(\mathbf{x}_{t},\mathbf{u}_{t})&\text{otherwise}\end{cases}$ (8)
Fig. 6 shows a study of using non-differentiable terminal penalty for both RL and differentiable simulation. The results show that adding a final penalty can greatly affect how well reinforcement learning (RL) works. Without a penalty, RL might get trapped in a local minimum. However, with a large penalty at the end, RL can achieve better task reward as well as more robust control performance. This is because RL optimizes a discounted return, which estimates “how good” it is to be in a given state. RL uses a state-value function to encode this information
$V_{\pi}(s)=\mathbb{E}[G|S_{0}=s]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|S_{0}=s\right],$ (9)
On the other hand, a terminal penalty has no impact on differentiable simulation since the gradient of a constant value is equal to zero. As a result, differentiable simulation requires well-defined continuous functions, e.g., a potential function or control barrier functions for robust control.
Figure 6: A comparison of non-differentiable terminal penalty for policy training. Using a non-differentiable terminal penalty, PPO can achieve robust control performance, e.g., longer episode length. We use 1024 robots for simulation.
V-D Real World Experiment
Figure 7: Real-world deployment using a Mini Cheetah. All the experiments are conducted using one single blind policy.
We demonstrate the performance of our policy in the real world using Mini Cheetah. Fig. 7 shows several snapshots of the robot’s behavior using different gait patterns. We trained a blind policy in simulation using 64 robots and then transfer the policy directly to the real world without fine-tuning. The robot can walk forward and backward with different gait patterns and frequencies. Moreover, the policy proved robust, enabling the robot to manage certain disruptions, such as unexpected forces applied to its body and locomotion on deformable terrain.
VI Limitations
A fundamental advantage of reinforcement learning (zero-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32]. Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
Our system design requires the specification of foot position using heuristic methods, which is a limitation compared to RL. RL’s flexibility allows for more robust performance enhancements, such as implementing a simple termination penalty to discourage falling behaviors or employing constant survival rewards to promote longevity.
VII Conclusion
In conclusion, we propose a novel framework leveraging the advantages of differentiable simulators for the complex task of quadruped locomotion in the real world. Our framework overcomes the challenges of contact-rich environments for differentiable simulators by integrating a simplified rigid-body model with a non-differentiable model for whole-body dynamics. In this setting, the learned locomotion policy is updated by directly backpropagating the differentiable loss function through the simplified model while the state of the robot is aligned using the more precise whole-body dynamics model. This framework not only introduces the advantages of faster and more stable training convergence of differential simulators to legged robots but also ensures the accurate simulation of complex contact forces. Notably, our experiments demonstrate that the proposed framework can train a policy to walk in minutes using only one simulation environment, a setting in which model-free RL falls short. Furthermore, the policy, aligned with an accurate simulator, seamlessly transfers to the real world in a zero-shot fashion.
An interesting future application of our simulator alignment is the task of learning directly on the hardware, where the locomotion policy is continually updated using the differentiable model in an online fashion while aligning robot states directly with real-world states. Beyond this, the proven training stability and fast training convergence of our differential simulator framework holds the promise for addressing the high-sample inefficiency in tasks like vision-based policy for navigation. In summary, our work bridges the gap between differentiable simulators and real-world legged locomotion, opening avenues for numerous future applications.
VIII Acknowledgments
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT). The authors especially thank Yuang Zhang for his help with the software development and Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control. Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.

### v2

@@
 The policy gradient can generally be expressed as the following
-$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R(\bm{\tau})],$
-where it stands as a zero-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$.
+$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R_{t}(\bm{\tau})],$
+where it stands as a zeroth-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$.
+Here, $R_{t}(\bm{\tau})$ represents the sum of future rewards.
 Despite impressive real-world performance, this approach to gradient estimation is known for its high variance.
@@
 However, the success of this approach heavily depends on the quality of expert demos and lacks generalizability and adaptability to new tasks.
-In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zero-order counterparts and hold great potential for more stable training and faster convergence.
-Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zero-order methods.
+In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence.
+Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zeroth-order methods.
 Figure 1: Graphical model for policy learning using differentiable simulation, in which both the loss function and the dynamics are differentiable.
@@
 Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes.
-Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zero-order methods, particularly model-free policy gradients.
+Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zeroth-order methods, particularly model-free policy gradients.
 We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
@@
 Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27].
-In principle, policy training through differentiable simulation allows better convergence by replacing the zero-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33].
+In principle, policy training through differentiable simulation allows better convergence by replacing the zeroth-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33].
 However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
@@
 Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$:
-$\mathbf{f}=\mathbf{J}\bm{\tau}.$
+$\mathbf{f}=(\mathbf{J}^{T})^{-1}\bm{\tau}.$
 The continuous nature of PD control enables the backpropagation of policy gradients. Consequently, our simulation framework treats the PD controller as an explicit, differentiable layer.
@@
 Following [23], the policy gradient can be expressed as follows
-$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{i}}{\partial x_{i}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
+$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
 where the matrix of partial derivatives $\partial x_{j}/\partial x_{j-1}$ is the Jacobian of the dynamical system $f$.
@@
 VI Limitations
-A fundamental advantage of reinforcement learning (zero-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32].
+A fundamental advantage of reinforcement learning (zeroth-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32].
 Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
@@
 864042 (AGILEFLIGHT).
-The authors especially thank Yuang Zhang for his help with the software development and Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control.
+The authors especially thank Yuang Zhang for his help with the software development, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for the valuable feedback.
 Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.

### v3

@@
 Abstract
-While most recent advancements in legged robot control have been driven by model-free reinforcement learning, we explore the potential of differentiable simulation.
-Differentiable simulation promises faster convergence and more stable training by computing low-variant first-order gradients using the robot model, but so far, its use for legged robot control has remained limited to simulation.
-The main challenge with differentiable simulation lies in the complex optimization landscape of robotic tasks due to discontinuities in contact-rich environments, e.g., quadruped locomotion.
-This work proposes a new, differentiable simulation framework to overcome these challenges.
-The key idea involves decoupling the complex whole-body simulation, which may exhibit discontinuities due to contact, into two separate continuous domains.
-Subsequently, we align the robot state resulting from the simplified model with a more precise, non-differentiable simulator to maintain sufficient simulation accuracy.
-Our framework enables learning quadruped walking in minutes using a single simulated robot without any parallelization.
-When augmented with GPU parallelization, our approach allows the quadruped robot to master diverse locomotion skills, including trot, pace, bound, and gallop, on challenging terrains in minutes.
-Additionally, our policy achieves robust locomotion performance in the real world zero-shot.
-To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot.
-This work provides several important insights into using differentiable simulations for legged locomotion in the real world.
-I Introduction
-Many grazing animals, such as giraffes, lambs, and calves, can learn to stand and walk within minutes to hours after birth.
-On the other hand, most machine learning algorithms, particularly traditional model-free reinforcement learning (RL), need massive parallelization to achieve a stable walking policy.
-How can legged robots master walking in minutes with a few trials and errors?
-Recent progress in legged robot control has largely been driven by GPU-accelerated massive simulation and reinforcement learning [15, 20, 24, 30, 22, 9, 39, 19].
-Central to these advancements is the policy gradient algorithm [35, 34, 6].
-The policy gradient can generally be expressed as the following
-$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R_{t}(\bm{\tau})],$
-where it stands as a zeroth-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$.
-Here, $R_{t}(\bm{\tau})$ represents the sum of future rewards.
-Despite impressive real-world performance, this approach to gradient estimation is known for its high variance.
-Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [31], increasing the number of samples, designing task-specific learning curriculums, employing distributed initialization techniques, and engineering suitable reward functions.
-As a result, RL struggles with complex tasks where data generalization is slow, like vision-based navigation.
-Despite the need for more effective learning algorithms, progress has been limited.
-Instead, research leans towards imitation learning, utilizing existing datasets or expert demonstrations for more consistent learning outcomes.
-Notably, the learning-by-cheating approach [4] is a popular imitation learning framework: first, you train an RL teacher with privileged information and then, you derive a student policy based on visual inputs.
-However, the success of this approach heavily depends on the quality of expert demos and lacks generalizability and adaptability to new tasks.
-In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence.
+This work explores the potential of using differentiable simulation for learning quadruped locomotion.
+Differentiable simulation promises fast convergence and stable training by computing low-variance first-order gradients using robot dynamics.
+However, its usage for legged robots is still limited to simulation.
+The main challenge lies in the complex optimization landscape of robotic tasks due to discontinuous dynamics.
+This work proposes a new differentiable simulation framework to overcome these challenges.
+Our approach combines a high-fidelity, non-differentiable simulator for forward dynamics with a simplified surrogate model for gradient back-propagation.
+This approach maintains simulation accuracy by aligning the robot states from the surrogate model with those of the precise, non-differentiable simulator.
+Our framework enables learning quadruped walking in simulation in minutes without parallelization.
+When augmented with GPU parallelization, our approach allows the quadruped robot to master diverse locomotion skills on challenging terrains in minutes.
+We demonstrate that differentiable simulation outperforms a reinforcement learning algorithm (PPO) by achieving significantly better sample efficiency while maintaining its effectiveness in handling large-scale environments.
+Our method represents one of the first successful applications of differentiable simulation to real-world quadruped locomotion, offering a compelling alternative to traditional RL methods.
+Video: https://youtu.be/weNq_w715xM
+Keywords: Differentiable Simulation, Legged Locomotion
+1 Introduction
+Figure 1: Graphical model for policy learning using differentiable simulation.
+Traditional model-free reinforcement learning (RL) often requires extensive parallelization to achieve a stable walking policy for legged robots [1].
+How can legged robots master walking quickly with a few trials and errors?
+Recent progress in legged robot control has been largely driven by combining model-free RL with massively parallelized simulation [1, 2, 3, 4, 5, 6, 7].
+Central to these advancements is the policy gradient algorithm [8, 9], where the gradient of a control policy can be calculated via $\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\nabla_{\theta}\log p(\tau;\theta)]$; which is a zeroth-order estimate of the true gradient of $J$ based on sampled trajectories following the policy $\pi_{\theta}$.
+Here, $R(\tau)$ is the reward of a trajectory $\tau$.
+Such zeroth-order gradient estimate provides a powerful and general framework for robot control since it can handle non-differentiable optimization objectives and discontinuous dynamics.
+However, the gradient has high variances.
+Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [10] and increasing the sample size.
+In robotics, leveraging well-established knowledges about robot dynamics can enable the construction of first-order gradient estimates $\nabla_{\theta}J(\theta)=\nabla_{\theta}R(\tau)$, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence.
 Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [11, 12, 13, 14, 15]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zeroth-order methods.
-Figure 1: Graphical model for policy learning using differentiable simulation, in which both the loss function and the dynamics are differentiable.
-However, the optimization landscape in robotic tasks is inherently complex, characterized by nonlinearities, non-smooth and discontinuous dynamics, non-convexities, and long temporal horizons.
-Additionally, deploying real-world robots introduces additional challenges, such as the sim-to-real gap, noisy state estimation, and system delay.
-Consequently, policy training through differentiable simulation can suffer from issues like local minimum and gradient vanishing/explosion.
-While differentiable simulation offers a promising avenue, its effectiveness in significantly accelerating policy learning in contact-rich environments and its applicability to real-world scenarios like dynamic quadruped control remains an open challenge in robotics.
-This work set out to explore the potential of differentiable simulation in accelerating policy training within contact-rich scenarios like dynamic legged locomotion.
-Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes.
-Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zeroth-order methods, particularly model-free policy gradients.
-We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
-Contribution: We show that differentiable simulation offers considerable advantages over model-free reinforcement learning for policy training in legged locomotion.
-Notably, we demonstrate that a single robot, without the need for parallelization, can quickly learn to walk within minutes using our approach.
+Differentiable simulation combines the advantages of model-based control and data-driven learning.
+Similar to numerical optimal control methods, such as shooting techniques, differentiable simulation utilizes robot dynamics to simulate trajectories and iteratively optimize decision variables using analytical gradients derived from the robot model.
+Like reinforcement learning, differentiable simulation benefits from large-scale simulations and deep learning to train neural network control policies that map observations directly to control actions.
+Despite its advantages, differentiable simulation faces several challenges, including discontinuous dynamics, complex optimization landscapes, and issues like exploding or vanishing gradients in long-horizon tasks, which can limit the effectiveness of first-order gradient methods.
+This work investigates the potential of training policies through surrogate models within differentiable simulation, specifically focusing on quadruped locomotion.
+We show that differentiable simulation offers considerable advantages over model-free RL for policy training in legged locomotion.
+Notably, we demonstrate that a single robot, without the need for parallelization can quickly learn to walk within minutes in simulation using our approach.
 Leveraging the advantage of GPUs for parallelized simulation, our robot learns diverse walking skills over challenging terrain in minutes.
 Specifically, we train a quadruped robot to walk with different gait patterns, including trot, pace, bound, and gallop, and with varying gait frequencies.
-While model-free RL can also achieve similar performance given sufficient parallelization, our approach requires much less data and yields better performance.
+While model-free RL can also achieve similar performance given sufficient parallelization, our approach requires much less data, achieving significantly better sample efficiency while maintaining its effectiveness in handling large-scale environments.
 More importantly, we show that the policy trained via differentiable simulation can be transferred to the real world directly without fine-tuning.
-To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot.
-This highlights the effectiveness of differentiable simulation for real-world applications in dynamic legged locomotion scenarios.
-The key to our approach is a novel policy training framework that combines the smooth gradients obtained from a simplified dynamics model for efficient backpropagation with the high fidelity of a more complex, non-differentiable simulator for accurate forward simulation.
-1) Decoupling Simulation Spaces: Rather than handling the whole body dynamics as a unified differentiable simulator, which may encounter non-smoothness due to contact interactions.
-We proposed to separate the simulation into two distinct spaces: the floating base space and the joint space.
-For the robot body, we employ an approximation using single rigid-body dynamics, which offers a continuous and effective representation of the floating base (the robot body).
-2) PD Control as A Differentiable Layer: The simulation within the joint space benefits from the inherent differentiability and smoothness of a Proportional-Derivative (PD) controller.
-This allows us to treat the PD control as an explicit, differentiable layer within our simulation framework.
-3) Alignment with Non-Differentiable Simulators: To address potential inaccuracies arising from our simplified rigid-body dynamics model, we incorporate a more precise, non-differentiable simulator.
-This non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our training pipeline remains grounded in realistic dynamics.
-II Related Work
-II-A Model Predictive Control
-Model predictive control (MPC) has a long history in legged locomotion [18, 16].
-MPC relies on fast online optimization and an accurate dynamics model to control the robot.
-To trade-off between model accuracy and computation speed, many methods use a reduced optimization model, such as single rigid-body dynamics [7, 11, 2].
-Research has also studied control using more complex formulations, including centroidal [10] and whole-body dynamics [26].
-Despite their success, MPC lacks the flexibility in its optimization objective design and robustness when facing disturbances or unexpected scenarios.
-Empirical evidence suggests that MPC tends to underperform when compared to data-driven approaches such as deep reinforcement learning.
-II-B Reinforcement Learning
-Deep reinforcement learning has emerged as a dominant approach for developing control policies in legged locomotion [15, 20, 24, 19].
-Recently, an important advancement in this area has been the introduction of IsaacGym [21], which is a GPU-accelerated simulator.
-IsaacGym can dramatically reduce the time required for dynamics simulation and policy training, enabling robots to learn to walk on flat ground in minutes [30].
-This advancement has significantly accelerated the pace of research in legged locomotion, opening new avenues to address increasingly complex challenges [38, 1, 9, 22, 39].
-Nevertheless, it is important to highlight that the progress in applying reinforcement learning to legged locomotion is primarily driven by the enhanced computational capabilities provided by modern GPUs rather than substantial breakthroughs in the underlying algorithms.
-Consequently, in scenarios where data collection cannot be accelerated through computational means, researchers may resort to alternative strategies such as “learning-by-cheating” to circumvent these limitations.
-This highlights a gap in the current approach to deep RL in legged locomotion, suggesting a need for innovative algorithmic developments that do not solely rely on computational advancements.
-II-C Differentiable Simulation
-Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27].
-In principle, policy training through differentiable simulation allows better convergence by replacing the zeroth-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33].
-However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
-Several approaches have been developed to tackle these challenges.
-For instance, SHAC [36] addresses the gradient issue by truncating trajectories into several smaller segments, which helps manage exploding or vanishing gradients.
-On the other hand, PODS [25] takes advantage of second-order gradients.
-This approach facilitates monotonic policy improvement and promises faster convergence than first-order methods.
-However, PODS’ effectiveness relies on very accurate computation of second-order derivatives from the differentiable simulator, making it sensitive to the precision of these derivatives.
-Another notable example is DiffMimic [29], which focuses on mimicking motion trajectories for physically simulated characters.
-Research in differentiable simulation has mainly been focused on speeding up policy optimization using first-order gradients, alleviating local minima problems, addressing gradient vanishing/explosion issues, or developing control policies for simulated characters.
-They show that differentiable simulation is a promising direction for improving sample efficiency.
-However, the capability of differentiable simulation to substantially speed up policy learning in contact-rich scenarios with physical interactions, as well as their practical applicability to real-world situations, such as quadruped locomotion, continues to be a significant challenge in robotics.
-Our research distinguishes itself from earlier studies by presenting a novel perspective that leads to an effective solution tailored for real-world robotic applications.
-III Methodology
-Figure 2: System overview of legged robot control using differentiable simulation.
-Our approach leverages the differentiability and smoothness of a single rigid-body dynamics for dynamics simulation.
-Additionally, we treat PD control as a differential layer in our computation graph.
-Finally, we use the state from a more accurate, non-differentiable simulator (IsaacGym) to align the simulated state in the simple rigid-body simulation.
-III-A Overview
-Fig.
-2 shows our training setup that employs differentiable simulation.
-The fundamental idea is that the main body of a legged robot can be modeled using differentiable and smooth single rigid-body dynamics, while the simulation of the robot’s four legs can be approached separately from its main body.
-We use a Proportional-Derivative (PD) controller to connect the legs dynamics with the main body.
-In the forward pass, this PD controller converts joint angles into torques, which then produce the contact forces needed for the robot’s rigid-body simulation.
-During backpropagation, the PD controller functions as a differentiable component, enabling gradients to be propagated back through the system.
-Finally, we use non-differentiable whole-body dynamics simulation to align the state produced by our simplified rigid-body dynamics.
-III-B Problem Formulation
+This work presents one of the earliest demonstrations of using differentiable simulation to train control policies for real quadruped robots, highlighting the potential of differentiable simulation for real-world applications.
+Contribution: The key to our approach is a novel policy training framework that combines the smooth gradients obtained from a simple surrogate dynamics model for efficient backpropagation with the high fidelity of a more complex, non-differentiable simulator for accurate forward simulation.
+Instead of simulating the quadruped robot using whole-body dynamics, which by definition has discontinuities due to contact, we propose separating the simulation into its floating base space and its joint space.
+For the robot body, we employ an approximation using single rigid-body dynamics, which offers a continuous and effective representation of the robot body.
+To address simulation errors arising from our simplified rigid-body dynamics model, we incorporate a more precise, non-differentiable simulator.
+This non-differentiable simulator can simulate complex contact dynamics and is used to align the state in our simplified model, thereby ensuring that our training pipeline remains grounded in realistic dynamics.
+Figure 2 provides an overview of our system.
+2 Methodology
+Figure 2: System overview of learning quadruped locomotion using differentiable simulation.
+Our approach decouples the robot dynamics into two separate spaces: joint and floating base spaces.
+We leverage the differentiability and smoothness of a single rigid-body dynamics for the robot’s main body, which takes the ground reaction force from its legs as the control inputs.
+We use the state from a non-differentiable simulator (IsaacGym) to align the state in the differentiable simulation.
+2.1 Problem Formulation
 We formulate legged robot control as an optimization problem.
@@
 An observation $o_{k}\in\mathcal{O}$ is generated at each time step based on the current state $x_{k}$ through a sensor model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{k}=h(x_{k})$.
-The system’s dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system as $x_{k+1}=x_{k}+\Delta t\cdot f(x_{k},u_{k})$.
-The discrete-time instants are $t_{k}=\Delta t\cdot k$, where $k$ ranges from 0 to $N$, thereby establishing a finite time horizon for the control problem.
+The system’s dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system as $x_{k+1}=f(x_{k},u_{k})$.
 At each time step $k$, the robot receives a cost signal $l_{k}=l(x_{k},u_{k})$, which is a function of the current state $x_{k}$ and the control input $u_{k}$.
@@
 The neural network takes the observation $o_{k}$ as input and outputs the control input $u_{k}$.
-The optimization objective is to find the optimal policy parameters $\theta^{\ast}$ by minimizing the total loss via gradient descent:
+The optimization objective is to find the optimal policy parameters $\theta^{\ast}$ by minimizing the total loss via gradient descent
 $\min_{\theta}$ $\mathcal{L}_{\theta}=\sum_{k=0}^{N-1}l(x_{k},u_{k})=\sum_{k=0}^{N-1}l(x_{k},\pi_{\theta}(o_{k}))$ (1)
-$\theta$ $\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta}$ (2)
+$\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta},$ (2)
 where $\alpha$ is the learning rate and $l(x_{k},u_{k})$ is the differentiable loss at simulation time step $k$.
-III-C Forward Simulation
-Legged robots are characterized by complex dynamics resulting from their multi-body configuration and the contact between their body and the environment.
-The contacts introduce discontinuities in the dynamics, making first-order gradient-based optimization algorithms difficult.
-An important aspect of our approach involves separating the simulation of the robot’s body from that of its joints.
-This design choice allows a simple implementation of the differentiable simulator.
-First, we represent the robot’s main body using single rigid-body dynamics.
-Single rigid-body dynamics have been shown to be useful for dynamic locomotion using online gradient-based optimization [7, 11].
+2.2 Forward Simulation
+We represent the robot’s main body using single rigid-body dynamics.
+Single rigid-body dynamics have been shown to be useful for dynamic locomotion using model predictive control [16, 17].
 The single rigid-body model offers a continuous representation of the robot base dynamics, avoiding the complex optimization landscape introduced by contacts.
@@
 $\dot{\mathbf{p}}_{WB}$ $=\mathbf{v}_{WB}$ $\dot{\mathbf{v}}_{WB}$ $=\frac{1}{m}\sum_{i}\mathbf{f}_{i}+\mathbf{g}$
-$\dot{\mathbf{q}}_{WB}$ $=\frac{1}{2}\Lambda(\bm{\omega}_{B})\cdot\mathbf{q}_{WB}$ $\dot{\bm{\omega}}_{B}$ $=\mathbf{I}^{-1}\left(\bm{\eta}-\bm{\omega}_{B}\times(\mathbf{I}\bm{\omega}_{B})\right)$
-In this approximation, the state of our system is
-$\mathbf{x}=[\mathbf{p},\mathbf{q},\mathbf{v},\bm{\omega}],$
-where $\mathbf{p}_{\text{WB}}\in\mathbb{R}^{3}$ is the position and $\mathbf{v}_{\text{WB}}\in\mathbb{R}^{3}$ is the linear velocity of the center of mass in the world frame ${W}$, We use a unit quaternion $\mathbf{q}_{\text{WB}}$ to represent the orientation of the body in the world frame and use $\bm{\omega}_{\text{B}}$ to denote the body rates in the body frame ${B}$.
-Here, $\mathbf{\Lambda}(\bm{\omega}_{B})$ is a skew-symmetric matrix.
-The control inputs are the ground reaction force $\mathbf{f}_{i}$ from the legs.
-To simulate the rigid-body dynamics, it is important to know the ground reaction force.
+$\dot{\mathbf{q}}_{WB}$ $=\frac{1}{2}\Lambda(\bm{\omega}_{B})\cdot\mathbf{q}_{WB}$ $\dot{\bm{\omega}}_{B}$ $=\mathbf{I}^{-1}\left(\bm{\eta}-\bm{\omega}_{B}\times(\mathbf{I}\bm{\omega}_{B})\right).$
+In this approximation, the state of our system is $\mathbf{x}=[\mathbf{p},\mathbf{q},\mathbf{v},\bm{\omega}],$ where $\mathbf{p}_{\text{WB}}\in\mathbb{R}^{3}$ is the position and $\mathbf{v}_{\text{WB}}\in\mathbb{R}^{3}$ is the linear velocity of the center of mass in the world frame ${W}$, We use a unit quaternion $\mathbf{q}_{\text{WB}}$ to represent the orientation of the body in the world frame and use $\bm{\omega}_{\text{B}}$ to denote the body rates in the body frame ${B}$.
+Here, $\mathbf{I}$ is the robot’s inertia tensor, $\bm{\eta}$ is the body torque, and $\mathbf{\Lambda}(\bm{\omega}_{B})$ is a skew-symmetric matrix.
+The control inputs are the ground reaction force $\mathbf{f}_{i}$ from the legs that have contacts.
+The ground reaction forces are required to simulate the rigid-body dynamics.
 One option for the control policy design is to output the ground reaction force directly, similar to the MPC design [16, 17].
 Another option is to control the robot in the joint space, e.g., output desired joint position, which allows more control authority for the policy and adaptive behavior.
-In this case, it is required to convert the neural network output (joint position) into the control input of the single rigid-body model, which is the ground reaction force.
-This conversion is achieved using a PD controller for forward propagation,
-$\bm{\tau}=\mathbf{k_{p}(q^{\text{ref}}-q)}+\mathbf{k_{d}(\dot{q}^{\text{ref}}-\dot{q})}$ (3)
+Our policy maps observations to the desired joint position $\mathbf{q}^{\text{ref}}$, while assuming zero joint velocity, i.e., $\mathbf{\dot{q}}^{\text{ref}}=\mathbf{0}$.
+In this case, the neural network output (joint position) is required to be converted into the control input of the single rigid-body model, which is the ground reaction force.
+This conversion is achieved using a PD controller for forward propagation
+$\bm{\tau}=\mathbf{k_{p}(q^{\text{ref}}-q)}+\mathbf{k_{d}(\dot{q}^{\text{ref}}-\dot{q})},$ (3)
 which calculates the required motor torque $\bm{\tau}$.
 Here, $\mathbf{k_{p}}$ and $\mathbf{k_{d}}$ are fixed gains, $\mathbf{q^{\text{ref}}}$ and $\mathbf{q}$ are the reference joint position and the current joint position separately, and $\mathbf{\dot{q}^{\text{ref}}}$ and $\mathbf{\dot{q}}$ are the reference joint velocity and the current joint velocity respectively.
-Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$:
-$\mathbf{f}=(\mathbf{J}^{T})^{-1}\bm{\tau}.$
-The continuous nature of PD control enables the backpropagation of policy gradients. Consequently, our simulation framework treats the PD controller as an explicit, differentiable layer.
-III-D Backpropagation Through Time
+Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$: $\mathbf{f}=(\mathbf{J}^{T})^{-1}\bm{\tau}.$ The continuous nature of PD control enables the backpropagation of policy gradients.
+Consequently, our simulation framework treats the PD controller as a differentiable layer.
+2.3 Backpropagation Through Time
 In differentiable simulation for policy learning, the backward pass is crucial for computing the analytic gradient of the objective function with respect to the policy parameters.
 Following [18], the policy gradient can be expressed as follows
-$\frac{\partial\mathcal{L}_{\theta}}{\partial\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
+$\nabla_{\theta}\mathcal{L}_{\theta}$ $=\frac{1}{N}\sum_{k=0}^{N-1}\left(\sum_{i=1}^{k}\frac{\partial l_{k}}{\partial x_{k}}\prod_{j=i}^{k}\underbrace{\left(\frac{\partial x_{j}}{\partial x_{j-1}}\right)}_{\text{differentiable dynamics}}\frac{\partial x_{i}}{\partial\theta}+\frac{\partial l_{k}}{\partial u_{k}}\frac{\partial u_{k}}{\partial\theta}\right),$ (4)
 where the matrix of partial derivatives $\partial x_{j}/\partial x_{j-1}$ is the Jacobian of the dynamical system $f$.
 Therefore, we can compute the policy gradient directly by backpropagating through the differentiable physics model and a loss function $l_{k}$ that is differentiable with respect to the system state and control inputs.
-A graphical model for gradient backpropagation in policy learning using differentiable simulation is given in Fig.
-1.
+A graphical model for gradient backpropagation in policy learning using differentiable simulation is given in Figure 1.
 Due to the usage of multiplication $\prod$, there are two potential issues in using Eq.
@@
 We tackle these two problems via short-horizon policy training.
-III-E Short-Horizon Policy Training
+2.4 State Alignment with A Non-Differentiable Simulator
+Due to the simplification of our single rigid-body dynamics and the decoupling of the simulation space, the robot state can diverge from its actual one.
+Over time, even minor discrepancies can accumulate, leading to unrealistic states and ultimately causing the failure of policy training.
+We proposed to align the body state in our differentiable simulation using information from other simulators that use accurate whole-body dynamics.
+Specifically, we align the robot state in our differentiable simulation with the state information from IsaacGym [19].
+We use the following equation to align the robot state $\mathbf{\hat{x}}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*(\mathbf{x}^{\text{diff}}_{t+1}-\mathbf{x}^{\text{diff, detach}}_{t+1}).$ Here, $\mathbf{x}^{\text{diff, detach}}_{t+1}$ and $\mathbf{x}^{\text{diff}}_{t+1}$ represent the same robot state of our differentiable simulator at time step $t+1$; hence, they share the same value.
+The word detach indicates that $\mathbf{\hat{x}}^{\text{diff, detach}}_{t+1}$ is detached from the computational graph for automatic differentiation.
+Therefore, we can reset the robot state using the value from the non-differentiable simulation during forward simulation: $\mathbf{\hat{x}}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*\bm{0}$.
+During backpropagation, the gradient of the state at a given time $t$ is computed as follows $\partial\mathbf{\hat{x}}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}=\bm{0}+\alpha*\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}.$ Here, $0<\alpha\leq 1$ is used to decay the gradient.
+The non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our differentiable training pipeline remains grounded in realistic dynamics.
+At the same time, our approach benefits from the simplified differentiable simulator, which offers smooth gradients for backpropagation.
+Figure 2 shows the computational graph of the forward propagation and backpropagation using state alignment.
+2.5 Short-Horizon Policy Training
 Although smoothed physical models (e.g., single rigid-body dynamics) improve the local optimization landscape, the complexity of the optimization problem escalates significantly in long-horizon problems involving extensive concatenation of simulation steps.
 The situation further deteriorates when the actions within each step are interconnected through a nonlinear and nonconvex neural network control policy.
-The complexity of the resulting reward landscape leads simple gradient-based methods to become trapped in local optima quickly.
+The complexity of the resulting optimization landscape can make gradient-based methods to become trapped in local optima quickly.
 Instead of directly solving a long-horizon policy training problem, following [13, 1], where long-horizon simulation tasks are truncated into short-horizon simulation, we utilize short-horizon policy learning to achieve stable gradient backpropagation over a short horizon.
-III-F Alignment with A Non-Differentiable Simulator
-Due to the simplification of our single rigid-body dynamics, the robot state can diverge from its actual one, eventually leading to unrealistic states due to compounding errors.
-We proposed to align the body state in our differentiable simulation using information from other simulators that use accurate whole-body dynamics.
-Specifically, we align the robot state in our differentiable simulation with the state information from IsaacGym [21].
-We use the following equation to align the robot state.
-$\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*(\mathbf{x}^{\text{diff}}_{t+1}-\mathbf{x}^{\text{diff, detach}}_{t+1})$ (5)
-Here, $\mathbf{x}^{\text{diff, detach}}_{t+1}$ and $\mathbf{x}^{\text{diff}}_{t+1}$ represent the same robot state of our differentiable simulator at time step $t+1$; hence, they share the same value.
-The word detach indicates that $\mathbf{x}^{\text{diff, detach}}_{t+1}$ is detached from the computational graph for automatic differentiation.
-Therefore, we can reset the robot state using the value from the non-differentiable simulation during forward simulation: $\mathbf{x}^{\text{diff}}_{t+1}=\mathbf{x}^{\text{non-diff}}_{t+1}+\alpha*\bm{0}$.
-During backpropagation, the gradient of the state at a given time $t$ is computed as follows
-$\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}=\bm{0}+\alpha*\partial\mathbf{x}^{\text{diff}}_{t+1}/\partial\mathbf{x}^{\text{diff}}_{t}.$ (6)
-Here, $0<\alpha\leq 1$ is used to decay the gradient.
-The non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our differentiable training pipeline remains grounded in realistic dynamics.
-At the same time, our approach benefits from the simplified differentiable simulator, which offers smooth gradients for backpropagation.
-Fig.
-2 shows the computational graph of the forward propagation and backpropagation using alignment.
-III-G Differentiable Loss Function
+Mathematically, short-horizon policy learning involves truncating a long-horizon trajectory of length $N$ into several shorter segments to compute the policy gradient shown in Eq.
+(4).
+For instance, for a trajectory of length 240, we divide it into ten segments of $N=24$.
+However, it’s important to note that while a shorter horizon could simplify computation and facilitate optimization, it may also limit the model’s ability to anticipate future events, introducing bias into the gradient calculation.
+This bias could, in turn, affect the types of tasks the model can effectively learn.
+2.6 Differentiable Loss Function
 A fundamental difference between policy training using differentiable simulation and reinforcement learning is that the loss function has to be differentiable when using differentiable simulation.
-On the one hand, RL allows the direct optimization of non-differentiable rewards, such as binary rewards $0/1$.
-On the other hand, differentiable simulation requires a smooth differentiable cost function to provide learning signals for the desired control inputs.
-We formulate a loss function tailored for velocity tracking, where the main objective is to follow a specified velocity command, denoted as $\mathbf{v}^{\text{ref}}$.
+RL allows the direct optimization of non-differentiable rewards, such as binary rewards $0/1$.
+On the contrary, differentiable simulation requires a smooth differentiable cost function to provide learning signals for the desired control inputs.
+We formulate a differentiable loss function $l(\mathbf{x}_{t},\mathbf{u}_{t})$ tailored for velocity tracking, where the main objective is to follow a specified velocity command, denoted as $\mathbf{v}^{\text{ref}}$.
 Additionally, we maintain the robot’s body height, represented by $p_{z}$.
-To enhance the robustness of this system, we incorporate several regularization terms: one to mitigate large angular velocity, thus controlling the body rate; another to limit the output action, preventing large control actions; and a term aimed at stabilizing the robot’s orientation using projected gravity vector.
-$l(\mathbf{x}_{t},\mathbf{u}_{t})$ $=$ $b\|\mathbf{v}-\mathbf{v}^{\text{ref}}\|^{2}+a\|p_{z}-p_{z}^{\text{ref}}\|+c\|\bm{\omega}\|^{2}$ (7)
-$+d\|\mathbf{u}\|^{2}+f\|\mathbf{g}_{\text{proj}}\|+g\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2}$
-Finally, we use a foot position loss function to provide a learning signal for the swing legs.
-This loss is critical for the swing leg since it contains information about the motor position.
-The swing trajectory is computed by fitting a quadratic polynomial over the lift-off, mid-air, and landing position of each foot, where the lift-off position is the foot location at the beginning of the swing phase, the landing position is calculated using the Raibert Heuristics [28].
-IV Experimental Setup
+To enhance the robustness of this system, we incorporate several regularization terms: one to mitigate a large angular velocity, thus controlling the body rate $\bm{\omega}$; another to limit the output action $\mathbf{u}$, preventing large control actions; and a term aimed at stabilizing the robot’s orientation using projected gravity vector $\mathbf{g}_{\text{proj}}$.
+The loss function is defined as
+$l(\mathbf{x}_{t},\mathbf{u}_{t})$ $=$ $a_{1}\|\mathbf{v}-\mathbf{v}^{\text{ref}}\|^{2}+a_{2}\|p_{z}-p_{z}^{\text{ref}}\|+a_{3}\|\bm{\omega}\|^{2}$ (5)
+$+a_{4}\|\mathbf{u}\|^{2}+a_{5}\|\mathbf{g}_{\text{proj}}\|+a_{6}\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2},$
+where $\mathbf{p}_{\text{foot}}$ is the foot position.
+3 Experimental Results
+3.1 A Toy Example: Control of A Double Integrator
+Figure 3: Control of a double integrator using optimal control, reinforcement learning, and differentiable simulation.
+(left): Learning curves.
+(middle): Trajectories of different control policies.
+We initialize the system at the same states for all methods.
+(right): DiffSim achieves control commands close to optimal control.
+Inspired by [20], where the author studied the connection between reinforcement learning and optimal control using a toy double integrator example, we begin by examining the same toy example: control of a double integrator, which is a fundamental problem in system and control theory.
+The double integrator is a second-order control system that models simple point mass dynamics in one-dimensional space.
+The state variables are position $x$ and velocity $\dot{x}$, with control inputs $u=\ddot{x}$.
+The objective is to drive the system state from any initial point to the origin, $[x,\dot{x}]=[0,0]$.
+Therefore, this problem can be solved by the Linear quadratic regulator (LQR), which can provide an optimal solution.
+We train a 2-layer multiplayer perception using backpropagation through time via differentiable simulation and the proximal policy optimization (PPO) method [10].
+We show that the policy trained using differentiable simulation (DiffSim) achieves nearly optimal performance within a handful of training iterations and samples.
+In contrast, even when scaling the number of simulation environments to 1024 and training the policy with considerably more iterations, PPO fails to achieve the same level of control performance as LQR or DiffSim.
+This toy example suggests that scaling might be insufficient for PPO to achieve an optimal solution.
+3.2 Learning to Walk with One Robot
+Figure 4: Learning to walk with one simulated robot. We run 10 experiments with different random seeds. The plot is smoothed using a moving average.
+This section explores whether a robot can learn to walk without parallelization and with very limited data points at each training iteration.
+We design a simple velocity-tracking task where the robot is required to follow a constant velocity in the $x$-axis, e.g., $v_{x}=0.2~{}m/s$.
+At each training iteration, we simulate 24-time steps with one single robot.
+Hence, each training iteration contains 24 data points.
+We trained the policy for a total of 1000 iterations.
+The result is given by the learning curve in Figure 4.
+Despite very limited data, our policy successfully learns to walk after minutes of training.
+As a comparison, PPO failed to achieve useful locomotion skills under the same condition.
+By providing first-order gradients with low variance, differentiable simulations offer a fundamental advantage: they allow for more efficient and reliable optimization of the control policy.
+This is because the gradient provides a clear direction for updating parameters, and the low variance ensures that this direction is consistent and reliable, even when data is limited.
+3.3 Learning Diverse Walking Skills on Challenging Terrains
+Figure 5: Learning to walk on challenging terrains, reinforcement learning versus differentiable simulation.
+Differentiable simulation exhibits significant advantages over PPO in terms of sample efficiency and learning stability.
+After training in simulation, the policy can be transferred to the real world without fine-tuning.
+[…]
+As the number of robots increases, the performance of both algorithms improves.
+Notably, the performance improvement for RL is much more significant than our approach.
+This indicates that the zeroth-order gradient estimates used by reinforcement learning are generally inaccurate and require many more samples to achieve stable training.
+On the contrary, the first-order gradient estimates used by our differentiable simulation can have very stable and accurate gradients, even given very limited simulation samples.
+Figure 5 demonstrates diverse walking skills over challenging terrains using a blind policy trained via differentiable simulation.
+[…]
+Robots Final Reward Total Training Time [min]
+PPO Ours PPO Ours
+4 $-25.82\pm 2.16$ $\mathbf{-3.83\pm 3.01}$ $7.27\pm 0.12$ $7.28\pm 0.12$
+16 $-8.07\pm 1.30$ $\mathbf{-2.03\pm 0.69}$ $7.54\pm 0.06$ $7.50\pm 0.08$
+64 $-6.42\pm 0.29$ $\mathbf{-2.49\pm 0.37}$ $8.46\pm 0.16$ $8.23\pm 0.07$
+1024 $-2.80\pm 0.13$ $\mathbf{-2.69\pm 0.06}$ $11.44\pm 0.13$ $11.55\pm 0.09$
+Table 1: A comparison of final reward and total training time. We run both methods for the same number of training iterations and collect the same amount of samples.
+Additionally, we compare the total training wall-clock time and the final performance of the resulting policies.
+Differentiable simulation often involves the creation of a complex computational graph to facilitate gradient computation.
+As the simulation horizon extends, the length of the computational graph can increase proportionally, leading to a substantial increase in the total training time.
+This increase can diminish its benefits in situations where gathering a large number of samples is both cheap and fast.
+Therefore, we ask the question: Can differentiable simulation be effectively applied to large-scale settings, for example, thousands of environments?
+To ensure a fair comparison, we ran both PPO and our training framework for the same number of training iterations, namely 1000 iterations.
+We compare the total training time, which includes the time for simulating dynamics and updating policies.
+As shown in Table 1, our results indicate that differentiable simulation can be as time efficient as PPO in large-scale environments despite the requirement of backpropagation through time (BPTT)—a process known for its computational demands.
+However, notice that the training time to reach a certain reward is much lower with our method and requires less robot and data.
+For example, our method requires only roughly 64 robots and 200 training iterations to solve this task, while PPO still performs less after 1000 training iterations with 1024 robots.
+4 Related Work
+Previous efforts in leveraging differentiable simulation to robot control can be broadly categorized into two classes: improving the simulator itself and enhancing the associated optimization algorithms.
+The first kind has led to the development of several general-purpose differentiable simulators, including Brax [11], Nimble [22], NeuralSim [23], and Dojo [24].
+These simulators enable the integration of robot models directly into policy optimization, allowing neural network control policies to be trained using more mathematically precise analytical gradients.
+However, as noted in [18, 25], in scenarios where the underlying system exhibits chaotic behavior—where small perturbations in initial conditions lead to significantly divergent states, such as in robotic simulations involving contact dynamics—these gradients can become unmanageable, leading to divergence in optimization processes.
+Our work highlights the benefits of employing well-behaved proxy dynamics as an alternative to the true robot dynamics involving contact.
+Such a design choice has its roots in model predictive control, which often leverages a simplified model for the optimization to control complex robotic systems [16, 17, 26].
+Besides developing a better simulator, various methods have been proposed to improve policy training algorithms.
+For example, [25] proposes an $\alpha$-order gradient estimator that aims to combine the efficiency of first-order estimates with the robustness of zeroth-order methods.
+SHAC [13] addresses local minima issues by learning a smooth value function and mitigates vanishing/exploding gradients through truncated backpropagation.
+Gradient norm clipping strategies have also proven effective in managing the exploding gradients issue [27].
+Furthermore, since differentiating through chaotic dynamics is challenging, it is often useful to learn an approximation of these functions and use the approximation for gradient computation.
+This approach commonly has a connection with model-based reinforcement learning [28, 29, 30, 31, 32], which involves learning a predictive model of the environment, followed by training a controller.
+Similar to [13], our method leverages a truncated backpropagation to facilitate optimization.
+Our method could be improved further by combining other ideas, such as learning a critic function or clipping the gradients.
+5 Limitations
+A key advantage of model-free RL lies in its ability to directly optimize task-level rewards [33], especially non-differentiable rewards.
+In contrast, differential simulation depends on the backpropagation of information from its objective function, making it challenging to explore novel solutions guided by task objectives.
+Consequently, our system requires specifying foot positions using the Raibert heuristic [34], as it cannot explore foot placement motion through velocity tracking loss alone.
+In addition, RL’s flexibility allows for more robust performance enhancements, such as implementing simple termination penalties to discourage falling behaviors or employing constant survival rewards to encourage continued operation.
+In addition, differentiable simulations are often tailored to specific optimization tasks and might not generalize well across different problems.
+It requires specific domain knowledge in both building the simulation pipeline and formulating the problem.
+To generalize our approach to other tasks, additional engineering efforts are required.
+6 Conclusion
+This work proposes a new perspective on learning quadruped locomotion policies using differentiable simulation by integrating a simplified rigid-body model with a non-differentiable model for whole-body dynamics.
+Our key insight is that a neural network control policy can be optimized directly by backpropagating the differentiable loss function through a simplified model while simultaneously aligning the robot’s state using a more accurate whole-body dynamics model.
+Future work should address several limitations of the proposed framework, such as removing the assumption of gait scheduling and the Raibert heuristic for foot contact sequence and foot position planning.
+Additionally, more sim-to-real analysis is helpful in evaluating the benefits of using a simplified model compared to the true but chaotic gradients or to the policy trained via domain randomization.
+Finally, we should demonstrate the true advantages of stable training introduced by the first-order gradient by tackling more challenging tasks, such as large-scale training and vision-based control.
+7 Acknowledgment
+This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No.
+101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No.
+864042 (AGILEFLIGHT).
+The authors especially thank Yuang Zhang for his help with the differentiable simulation development and for the useful discussion, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for his valuable feedback.
+Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.
+Appendix
+Appendix 1 Double Integrator
+A double integrator system is characterized by its position and velocity.
+The control input $u$ directly controls the acceleration of the system.
+For a fair comparison, we formulate the problem as a discrete-time finite-horizon optimal control problem.
+The discrete-time state space representation of a double integrator is
+$x_{k+1}=Ax_{k}+Bu_{k}$
+where $A=\begin{bmatrix}1&d_{t}\\ 0&1\end{bmatrix}$ and $B=\begin{bmatrix}d_{t}^{2}/2\\ dt\end{bmatrix}$.
+Here, $d_{t}$ is the simulation time step.
+The objective is to minimize a quadratic loss function
+$J$ $=x_{N}^{T}Q_{f}x_{N}+\sum_{k=0}^{N-1}(x_{k}^{T}Qx_{k}+u_{k}^{T}Ru_{k})$
+$Q$ $=Q_{f}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},R=[1].$
+This is a discrete-time finite-horizon Linear–Quadratic Regulator problem, where the optimal control law can be found using dynamic programming
+$u_{k}^{\ast}$ $=-(R+B^{T}P_{k+1}B)^{-1}B^{T}P_{k+1}Ax_{k},\quad k=0,1,\cdots,N-1,$
+where $P_{N}=Q_{f}$ and $P_{k}$ can be found from the Riccati recursion.
+Appendix 2 Experimental Setup
 […]
@@
 $\mathbf{g}_{\text{projected}}$ 3
-TABLE I: Policy observation and action.
-[…]
-V Experimental Results
-V-A Learning to Walk with One Robot
-This section explores whether a robot can learn to walk in a single simulation environment without parallelization.
-We design a simple velocity tracking task where the robot is required to follow a constant velocity in the $x$-axis, e.g., $v_{x}=0.2~{}m/s$.
-At each training iteration, we simulate 24 times steps with one single robot.
-Hence, each training iteration contains only 24 data points.
-We trained the policy for a total 1000 iterations, which corresponded to 24,000 data points and 4 minutes of real-time experiences in total.
-The result is given by the learning curve in Fig.
-3.
-Surprisingly, despite extremely limited data, our policy successfully learns to walk after minutes of training.
-As a comparison, model-free reinforcement learning failed to yield positive outcomes under the same condition.
-Figure 3: Learning to walk with one simulated robot. We run 10 experiments with different random seeds. The plot is smoothed using a moving average.
-V-B Learning Diverse Walking Skills on Challenging Terrains
-[…]
-Fig.
-5 shows the learning curves using different numbers of robots for policy training.
-We compare the learning performance of our method with a model-free RL algorithm (PPO) [31].
-Given minimal samples, e.g., only four robots, RL has a plodding convergence speed and does not learn meaningful walking skills, e.g., the policy constantly falls after a few simulation steps.
-In contrast, differentiable simulation achieves much higher rewards and can acquire useful walking skills, albeit with relatively low success rates.
-As the number of robots increases, the performance for both algorithms improved.
-Notably, the performance improvement for RL is much more significant than our approach.
-This indicates the that the zero-order gradient estimates used by reinforcement learning is generally inaccurate and requires much more samples to achieve stable training.
-On the contrary, the first-order gradient estimates used by our differentiable simulation can have very stable and accurate gradients even given very limited simulation samples.
-Fig.
-5 demonstrates diverse walking skills over challenging terrains using a blind policy trained via differentiable simulation.
-Figure 4: Learning to walk on challenging terrains, reinforcement learning versus differentiable simulation.
-Differentiable simulation achieves faster and more stable convergence, given limited simulation samples.
-Additionally, we compare the training wall-clock time and the performance of the resulting policies between the two methodologies.
-Differentiable simulation often involves the creation of a complex computational graph to facilitate gradient computation, especially for backpropagation through time.
-As the simulation horizon extends, the length of the computational graph can increase proportionally, leading to a substantial increase in the total training time.
-This increase can diminish its benefits in situations where gathering a large number of samples is both cheap and fast.
-Robots Final Reward Training Time [min]
-PPO Ours PPO Ours
-4 $-25.82\pm 2.16$ $\mathbf{-3.83\pm 3.01}$ $\mathbf{7.27\pm 0.12}$ $7.28\pm 0.12$
-16 $-8.07\pm 1.30$ $\mathbf{-2.03\pm 0.69}$ $7.54\pm 0.06$ $\mathbf{7.50\pm 0.08}$
-64 $-6.42\pm 0.29$ $\mathbf{-2.49\pm 0.37}$ $8.46\pm 0.16$ $\mathbf{8.23\pm 0.07}$
-1024 $-2.80\pm 0.13$ $\mathbf{-2.69\pm 0.06}$ $\mathbf{11.44\pm 0.13}$ $11.55\pm 0.09$
-TABLE II: Learning to walk on challenging terrains, reinforcement learning versus differentiable control. We compare the final reward and the total training walk-clock time.
-Figure 5: Diverse locomotion skills over challenging terrains using a blind policy trained via differentiable simulation.
-As presented in Table II, a differentiable simulation achieves comparable overall training time.
-This efficiency is achieved by employing a truncated simulation timestep, which results in a smaller computational graph and enables fast and effective backpropagation.
-RL requires marginally longer training time due to its mini-batch and multi-epoch training scheme, in contrast, differentiable simulation only uses one gradient update.
-Differentiable simulation achieved overall higher reward during evaluation.
-V-C On the Importance of Non-differentiable Terminal Penalty
-This subsection highlights one important benefit of Reinforcement Learning (RL) compared to differentiable simulation: RL can significantly enhance its robustness by directly optimizing through non-differentiable rewards or penalties.
-Specifically, we use an non-differentiable value $p=200$ to penalize the robot (Eq.
-7) when the robot experiences termination during training, e.g., falling on the ground.
-$r(\mathbf{x}_{t},\mathbf{u}_{t})=\begin{cases}-l(\mathbf{x}_{t},\mathbf{u}_{t})-p&\text{if termination}\\ -l(\mathbf{x}_{t},\mathbf{u}_{t})&\text{otherwise}\end{cases}$ (8)
+Table 2: Policy observation and action for quadruped locomotion.
+Appendix 3 Foot Trajectory Planning
+We use the foot position loss $\|\mathbf{p}_{\text{foot}}-\mathbf{p}_{\text{foot}}^{\text{ref}}\|^{2}$ to provide a learning signal for the swing legs.
+This loss term is critical for the swing leg since it contains information about the motor position.
+Following prior works [35, 36], the swing leg trajectory can be computed by fitting a quadratic polynomial over the lift-off $\mathbf{p}^{\text{lift}}_{\text{foot}}$, mid-air $\mathbf{p}^{\text{air}}_{\text{foot}}$, and landing position $\mathbf{p}^{\text{land}}_{\text{foot}}$ of each foot, where the lift-off position is the foot location at the beginning of the swing phase, the landing position $\mathbf{p}^{\text{land}}_{\text{foot}}$ is calculated using the Raibert Heuristics [34], which is expressed as the following function
+$\mathbf{p}^{\text{land}}_{\text{foot}}=\mathbf{p}^{\text{hip}}_{\text{foot}}+\mathbf{v}^{\text{CoM}}T_{\text{stance}}/2$
+where $T_{\text{stance}}$ is the expected time the foot will spend on the ground, $\mathbf{p}^{\text{hip}}_{\text{foot}}$ is the location on the ground beneath the robot’s hip, $\mathbf{v}^{\text{CoM}}$ is the body velocity projected on the $xy$-plane.
+The desired contact state of each leg.
+e.g., swing or stance, is determined via a gait generator.
+The gait is modulated by a phase variable $\phi\in[0,2\pi]$.
+The phase is defined through a dynamic function $\phi_{t+1}=\phi_{t}+2\pi f\Delta t$, where $f$ is the stepping frequency.
+We can design different locomotion patterns by adapting the stepping frequency $f$ and the phase difference between each leg.
+Appendix 4 On the Importance of Non-differentiable Terminal Penalty
+We highlight one important benefit of RL compared to differentiable simulation: RL can significantly enhance its robustness by directly optimizing through non-differentiable rewards or penalties.
+Specifically, we use a non-differentiable value $p=200$ to penalize the robot when the robot experiences termination during training, e.g., falling on the ground or lifting its legs above its body.
+$r(\mathbf{x}_{t},\mathbf{u}_{t})=\begin{cases}-l(\mathbf{x}_{t},\mathbf{u}_{t})-p&\text{if termination}\\ -l(\mathbf{x}_{t},\mathbf{u}_{t})&\text{otherwise}.\end{cases}$
 Fig.
 6 shows a study of using non-differentiable terminal penalty for both RL and differentiable simulation.
-The results show that adding a final penalty can greatly affect how well reinforcement learning (RL) works.
+The results show that adding a final penalty can greatly affect how well RL works.
 Without a penalty, RL might get trapped in a local minimum.
-However, with a large penalty at the end, RL can achieve better task reward as well as more robust control performance.
+However, with a large penalty at the end, RL can achieve better task rewards as well as more robust control performance.
 This is because RL optimizes a discounted return, which estimates “how good” it is to be in a given state.
 RL uses a state-value function to encode this information
-$V_{\pi}(s)=\mathbb{E}[G|S_{0}=s]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|S_{0}=s\right],$ (9)
-On the other hand, a terminal penalty has no impact on differentiable simulation since the gradient of a constant value is equal to zero.
+$V_{\pi}(s)=\mathbb{E}[G|S_{0}=s]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|S_{0}=s\right].$
+On the other hand, a terminal penalty has no impact on differentiable simulation since the gradient of a constant value is equal to zero, and we do not leverage a state value function.
 As a result, differentiable simulation requires well-defined continuous functions, e.g., a potential function or control barrier functions for robust control.
@@
 We use 1024 robots for simulation.
-V-D Real World Experiment
-Figure 7: Real-world deployment using a Mini Cheetah. All the experiments are conducted using one single blind policy.
-We demonstrate the performance of our policy in the real world using Mini Cheetah.
-Fig.
-7 shows several snapshots of the robot’s behavior using different gait patterns.
-We trained a blind policy in simulation using 64 robots and then transfer the policy directly to the real world without fine-tuning.
-The robot can walk forward and backward with different gait patterns and frequencies.
-Moreover, the policy proved robust, enabling the robot to manage certain disruptions, such as unexpected forces applied to its body and locomotion on deformable terrain.
-VI Limitations
-A fundamental advantage of reinforcement learning (zeroth-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32].
-Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
-Our system design requires the specification of foot position using heuristic methods, which is a limitation compared to RL.
-RL’s flexibility allows for more robust performance enhancements, such as implementing a simple termination penalty to discourage falling behaviors or employing constant survival rewards to promote longevity.
-VII Conclusion
-In conclusion, we propose a novel framework leveraging the advantages of differentiable simulators for the complex task of quadruped locomotion in the real world.
-Our framework overcomes the challenges of contact-rich environments for differentiable simulators by integrating a simplified rigid-body model with a non-differentiable model for whole-body dynamics.
-In this setting, the learned locomotion policy is updated by directly backpropagating the differentiable loss function through the simplified model while the state of the robot is aligned using the more precise whole-body dynamics model.
-This framework not only introduces the advantages of faster and more stable training convergence of differential simulators to legged robots but also ensures the accurate simulation of complex contact forces.
-Notably, our experiments demonstrate that the proposed framework can train a policy to walk in minutes using only one simulation environment, a setting in which model-free RL falls short.
-Furthermore, the policy, aligned with an accurate simulator, seamlessly transfers to the real world in a zero-shot fashion.
-An interesting future application of our simulator alignment is the task of learning directly on the hardware, where the locomotion policy is continually updated using the differentiable model in an online fashion while aligning robot states directly with real-world states.
-Beyond this, the proven training stability and fast training convergence of our differential simulator framework holds the promise for addressing the high-sample inefficiency in tasks like vision-based policy for navigation.
-In summary, our work bridges the gap between differentiable simulators and real-world legged locomotion, opening avenues for numerous future applications.
-VIII Acknowledgments
-This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No.
-101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No.
-864042 (AGILEFLIGHT).
-The authors especially thank Yuang Zhang for his help with the software development, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for the valuable feedback.
-Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.
+Appendix 5 Ablation Study
+We conducted an ablation study to investigate the significance of the proposed state alignment mechanism.
+In our implementation, we use the state from IsaacGym to align the robot state within our simplified rigid-body differentiable simulation.
+To assess the impact of state alignment, we performed experiments where we removed the state alignment and compared the resulting learning curves to those obtained with state alignment.
+The results are presented in Figure 7.
+The results indicate that without state alignment, the robot fails to learn any useful walking skills.
+Figure 7: Ablation study for state alignment using IsaacGym.
+Appendix 6 Hyperparameters
+Parameter Value
+learning rate 0.001
+discount factor $\gamma$ 0.95
+GAE-$\lambda$ 0.95
+learning epoch 10
+policy network MLP [256, 256]
+value network MLP [256, 256]
+clip range 0.2
+entropy coefficient 0.002
+number of epoch 10
+Table 3: PPO hyperparameters.
+Parameter Value
+learning rate 0.001
+policy network MLP [256, 256]
+gradient decay factor $\alpha$ 0.9
+Table 4: Hyperparameters for policy training using differentiable simulation.